### v1

# Primitive Quantum Gates for an $SU(3)$ Discrete Subgroup: $\Sigma(36\times 3)$

Abstract
We construct the primitive gate set for the digital quantum simulation of the 108-element $\Sigma(36\times 3)$ group. This is the first time a nonabelian crystal-like subgroup of $SU(3)$ has been constructed for quantum simulation. The gauge link registers and necessary primitives – the inversion gate, the group multiplication gate, the trace gate, and the $\Sigma(36\times 3)$ Fourier transform – are presented for both an eight-qubit encoding and a heterogeneous three-qutrit plus two-qubit register. For the latter, a specialized compiler was developed for decomposing arbitrary unitaries onto this architecture.
preprint: FERMILAB-PUB-24-0132-SQMS-T, TUM-HEP-1506/24
I Introduction
Classical computers face significant challenges in simulating lattice gauge theories due to inherent exponentially large Hilbert spaces with the lattice volume. Monte Carlo simulations in Euclidean time are generally used to circumvent this problem. However, this approach also fails when we are interested in the real time dynamics of the system or in the properties of matter at finite density due to the sign problem [1, 2, 3, 4, 5, 6, 7, 8].
Quantum computers provide a natural way of simulating lattice gauge theories. Yet, they are currently limited to a small number of qubits and circuit depths. Gauge theories contain bosonic degrees of freedom and have a continuous symmetry, e.g. Quantum chromodynamics (QCD) with $SU(3)$ local symmetry. Storing a faithful matrix representation of $SU(3)$ to double precision would require $\mathcal{O}(10^{3})$ qubits per link – far beyond accessibility to near-term quantum computers. Moreover, these qubits being noisy significantly limits the circuit depths that can be reliably performed on these devices. Therefore, studying lattice gauge theories with current and near-future quantum computers requires efficient digitization methods of gauge fields as well as optimized computational subroutines. Finally, it is important to note that a choice of digitization method affects the computation cost.
To this end, several digitization methods have been proposed in the past decade to render the bosonic Hilbert space finite. Traditionally, most approaches have considered only qubit devices. However, due to recent demonstrations demonstration of qudit gates, there has been an increasing interest in qudit-based digitization methods [9, 10, 11, 12, 13, 14, 15, 16]. One digitization method utilizes the representation (electric field) basis, and impose a cut off on the maximal representation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 13, 32, 33, 34, 35, 36]. Another recent proposal [37, 16], the $q$-deformed formulation renders a finite dimension by replacing the continuous gauge group with a so-called quantum group [38]. Moreover, the loop-string-hadron formulation explicitly enforces gauge invariance on the Hilbert space [39, 40, 41, 42, 43] before truncation. Each of the above methods is extendable to the full gauge group, i.e. it has an infinite-dimensional limit. Other methods with begin with different formulations or perform different approximations exist such as light-front quantization [44, 45, 46], conformal truncation [47], strong-coupling and large-$N_{c}$ expansions [48, 49].
Another approach is to try and formulate a finite-dimensional Hilbert space theory with continuous local gauge symmetry which is in the same universality class as the original theory. For example, the author of Ref. [50] constructed a $SU(2)$ gauge theory where each link Hilbert space is five-dimensional. A generalization to $SU(N)$, however, was not obtained due to a spurious $U(1)$ symmetry. Later, different finite-dimensional formulations were found for $SU(2)$ gauge theories [51], the smallest of which being four-dimensional. Recently, a method inspired from non-commutative geometry was used to construct a $SU(2)$ gauge theory in 16-dimensional Hilbert space on each link as well as a generalization to $U(N)$ gauge theory [52]. Another finite-dimensional digitization known as quantum link models uses an ancillary dimension to store a quantum state [53]. This method can be extended to an arbitrary $SU(N)$, and has been further investigated in Refs. [54, 55, 56, 53, 57, 58, 59, 60, 61, 62, 63]. Although this approach may greatly simplify the cost of digitization, establishing the universality class is non-trivial [52, 51, 64].
Another promising approach to digitization is the discrete subgroup approximation [65, 66, 67, 68, 69, 70, 71, 72, 14, 73, 74, 75, 65, 66, 67, 68, 69, 70, 71, 72, 14, 73, 74, 76, 77, 78, 79, 39]. This method was explored early on in Euclidean lattice field theory to reduce computational resources in Monte Carlo simulations of gauge theories. Replacing $U(1)$ by $\mathbb{Z}_{N}$ was considered in [80, 81]. Extensions to the crystal-like subgroups of $SU(N)$ were made in Refs. [82, 83, 84, 66, 67, 69, 85, 86, 87], including with fermions [88, 89]. Theoretical studies revealed that the discrete subgroup approximation corresponds to continuous groups broken by a Higgs mechanism [90, 91, 92, 93, 94]. On the lattice, this causes the discrete subgroup to poorly approximate the continuous group below a freezeout lattice spacing $a_{f}$ (or beyond a coupling $\beta_{f}$), see Fig. 1.
The discrete group approximation has several significant advantages over many other methods discussed above. It is a finite mapping of group elements to integers that preserves a group structure; therefore it avoids any need for expensive fixed- or floating-point quantum arithmetic. The inherent discrete gauge structure further allows for coupling the gauge redundancy to quantum error correction [95, 77]. Additionally, while other method in principle need to increase both circuit depth and qubit count to improve the accuracy of the Hilbert space truncation, the discrete group approximation only needs to include additional terms into the Hamiltonian [96, 86]
Table 1: Parameters of a crystal-like subgroups of $SU(3)$. $\Delta S$ is the gap between $\mathbb{1}$ and the nearest neighbors to it. $\mathcal{N}$ is number of group elements that neighbor the $\mathbb{1}$.
$G$ $\Delta S$ $\mathcal{N}$ $\beta^{2+1d}_{f}$ $\beta^{3+1d}_{f}$
$\Sigma(36\times 3)$ $\frac{2}{3}$ 18 3.78(2) 2.52(3)
$\Sigma(72\times 3)$ $\frac{2}{3}$ 54 3.2(1)111from [82]
$\Sigma(216\times 3)$ $1-\frac{1}{3}\left(\cos\frac{\pi}{9}+\cos\frac{2\pi}{9}\right)$ 24 3.43(2)1
$\Sigma(360\times 3)$ $\frac{5-\sqrt{5}}{6}$ 72 3.935(5)222from [67]
In this work, we consider the smallest crystal-like subgroup of a $SU(3)$ with a $\mathbb{Z}_{3}$ center – $\Sigma(36\times 3)$ which has 108 elements. These elements can be naturally encoded into a register consider of 8 qubits or 3 qutrits & 2 qubits. A number of smaller nonabelian subgroups of $SU(2)$ have been considered previously: the $2N$-element dihedral groups $D_{N}$ [65, 97, 98, 99], the 8-element $\mathbb{Q}_{8}$ [14], the crystal-like 24-element $\mathbb{BT}$ [100], and the crystal-like 48 element $\mathbb{BO}$ [101]. From Fig. 1, we observe that freezeout occurs far before the scaling regime. This implies that the Kogut-Susskind Hamiltonian (which can be derived from the Wilson action) is insufficient for $\Sigma(36\times 3)$ to approximate $SU(3)$, but classical calculations suggest with modified or improved Hamiltonians $H_{I}$ may prove sufficient for some groups [67, 69, 85, 86].
Figure 1: Euclidean calculations of lattice energy density $\langle E_{0}\rangle$ of $\Sigma(36\times 3)$ as measured by the expectation value of the plaquette as a function of Wilson coupling $\beta$ on $8^{d}$ lattices for (top) $(2+1)$-dimensions (bottom) $(3+1)$-dimensions. The shaded region indicates $\beta\geq\beta_{s}$.
This paper is organized as follows. In Sec. II, the group theory needed for $\Sigma(36\times 3)$ are summarized and the digitization scheme is presented. Sec. IV demonstrates the quantum circuits for the four primitive gates required for implementing the group operations: the inversion gate, the multiplication gate, the trace gate, and the Fourier transform gate. Using these gates, Sec. V presents a resource estimates for simulating $3+1d$ $SU(2)$. We conclude and discuss future work in Sec. VI.
II Properties of $\Sigma(36\times 3)$
$\Sigma(36\times 3)$ is a discrete subgroup of SU(3) with 108 elements. The group elements, $g$, of $\Sigma(36\times 3)$ can be written in the following ordered product or otherwise known as strong generating set. That is, all the group elements, $g$, can be enumerated as a product of left or right transversals such that
$g=\omega^{p}C^{q}E^{r}V^{2s+t},$ (1)
where $0\leq p,~{}q,~{}r\leq 2$ and $0\leq s,~{}t\leq 1$. This indicates that either 8 qubits (2 each for $p$ , $q$, and $r$, and one each for $s$ and $t$) or 3 qutrits ($p$, $q$, and $r$) and 2 qubits ($s$ and $t$) will be required to store the group register. Because the indices $p,~{}q,$ and $r$ take on values between 0 and 2, there exists an ambiguity in mapping the three-level states to a pair of two-level systems. We use the mapping $|0\rangle_{3}=|00\rangle_{2}$, $|1\rangle_{3}=|01\rangle_{2}$, and $|2\rangle=|10\rangle_{2}$ with the $|11\rangle_{2}$ state being forbidden. Throughout this work, we will use $|\rangle_{3}$ to denote a three-level state and $|\rangle_{2}$ to denote a two-level state when there is the possibility of ambiguity. In this way, the index $p$ is decomposed in binary as $p=p_{0}+2p_{1}$ and encoded as the state $|p\rangle_{3}=|p_{1}p_{0}\rangle_{2}$. This process is done similarly for $q$ and $r$.
The strong generating set shown in Eq. 1 explicitly builds the presentation of the group from subgroups. In this way primitive gates for smaller discrete groups can be used as building blocks to construct efficient primitive gates of larger groups [102, 100, 101]. In the case of $\Sigma(36\times 3)$, the subgroups of interest are as follows: $\omega^{p}$ generates the subgroup $\mathbb{Z}_{3}$; $\omega^{p}C^{q}$ generates the subgroup $\mathbb{Z}_{3}\times\mathbb{Z}_{3}$; $\omega^{p}C^{q}E^{r}$ generates the subgroup $\Delta(27)$; $\omega^{p}C^{q}E^{r}V^{2s}$ generates $\Delta(54)$. Detailed information regarding these subgroups can be found in Ref. [103].
It is useful to have the irreducible representations, irreps, of $\Sigma(36\times 3)$ for deriving a quantum Fourier transformation (see Sec. IV). This group has 14 irreducible representations (irreps). There are four one-dimensional (1d) irreps, eight three-dimensional irreps (3d), and two four-dimensional (4d) irreps. The 1d irreps are:
$\rho^{(1)}_{a}(g)=i^{a(2s+t)},~{}0\leq a\leq 3.$ (2)
The eight 3d irreps can be written as
$\rho^{(3)}_{a,b}(g)=(-1)^{abt}\omega^{(1+b)p}C^{(1+b)q}E^{r}(i^{a}V)^{2s+(-1)^{b}t},$ (3)
where $0\leq a\leq 3$ and $0\leq b\leq 1$ and the matrices $\omega$, $C$, $E$, and $V$ are given by
$\omega=e^{2\pi i/3}$ $,~{}C=\text{Diag}(1,\omega,\omega^{2}),$
[…]
The irrep $\rho_{0,0}^{(3)}$ corresponds to the faithful irrep that resembles the fundamental irrep of $SU(3)$. The 4d irreps are given by Eq. (1) with
$\begin{split}\rho^{(4)}_{b}(\omega)=&\mathbb{1},~{}~{}\rho^{(4)}_{b}(C)=\text{Diag}(\omega^{b},\omega,\omega^{2b},\omega^{2}),\\ \rho^{(4)}_{b}(E)=&\text{Diag}(\omega,\omega^{2b},\omega^{2},\omega^{b}),\\ \rho^{(4)}_{b}(V)=&\begin{pmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ \end{pmatrix}.\end{split}$ (5)
In addition for conciseness we provide the character table from Ref [103] in Tab. 2, which will be useful in constructing the trace gate.
As we proceed with constructing primitive gates (see Sec. IV),the following “reordering” relations are useful:
$EC=$ $\omega CE,~{}VC=EV,~{}\text{and}~{}VE=C^{2}V.$ (6)
One can extend the relations above to derive the generalized reordering relations:
$V^{2s+t}C^{q}=$ $C^{(1+s)q(1-t)}E^{t(1+s)q}V^{2s+t},$
$E^{r}C^{q}=$ $\omega^{rq}C^{q}E^{r},$
$V^{2s+t}E^{r}=$ $C^{t(s+1)(3-r)}E^{(1-t)r(1+s)}V^{2s+t}.$ (7)
Table 2: Character table of $\Sigma(36\times 3)$ with $\omega=e^{2\pi i/3}$ [103].
Size 1 1 1 12 12 9 9 9 9 9 9 9 9 9
Ord. 1 3 3 3 3 2 6 6 4 12 12 4 12 12
$\mathbf{1}^{(0)}$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$\mathbf{1}^{(1)}$ $1$ $1$ $1$ $1$ $1$ -$\text{-}1$$\text{-}1$$\text{-}1$ $i$ $i$ $i$$\text{-}i$$\text{-}i$$\text{-}i$
$\mathbf{1}^{(2)}$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ -$\text{-}1$$\text{-}1$$\text{-}1$$\text{-}1$$\text{-}1$$\text{-}1$
$\mathbf{1}^{(3)}$ $1$ $1$ $1$ $1$ $1$ -$\text{-}1$$\text{-}1$$\text{-}1$$\text{-}i$$\text{-}i$$\text{-}i$ $i$ $i$ $i$
$\mathbf{3}^{(0)}$ $3$ $3\omega$ $3\omega^{2}$ 00 00 -$\text{-}1$$\text{-}\omega$$\text{-}\omega^{2}$ $1$ $\omega$ $\omega^{2}$ $1$ $\omega$ $\omega^{2}$
$\mathbf{3}^{(1)}$ $3$ $3\omega$ $3\omega^{2}$ 00 00 $1$ $\omega$ $\omega^{2}$ $i$ $i\omega$ $i\omega^{2}$$\text{-}i$$\text{-}i\omega$$\text{-}i\omega^{2}$
$\mathbf{3}^{(2)}$ $3$ $3\omega$ $3\omega^{2}$ 00 00 -$\text{-}1$$\text{-}\omega$$\text{-}\omega^{2}$$\text{-}1$$\text{-}\omega$$\text{-}\omega^{2}$$\text{-}1$$\text{-}\omega$$\text{-}\omega^{2}$
$\mathbf{3}^{(3)}$ $3$ $3\omega$ $3\omega^{2}$ 00 00 $1$ $\omega$ $\omega^{2}$$\text{-}i$$\text{-}i\omega$$\text{-}i\omega^{2}$ $i$ $i\omega$ $i\omega^{2}$
$\mathbf{3}^{(0)*}$ $3$ $3\omega^{2}$ $3\omega$ 00 00 -$\text{-}1$$\text{-}\omega^{2}$$\text{-}\omega$ $1$ $\omega^{2}$ $\omega$ $1$ $\omega^{2}$ $\omega$
$\mathbf{3}^{(1)^{*}}$ $3$ $3\omega^{2}$ $3\omega$ 00 00 $1$ $\omega^{2}$ $\omega$$\text{-}i$$\text{-}i\omega^{2}$$\text{-}i\omega$ $i$ $i\omega^{2}$ $i\omega$
$\mathbf{3}^{(2)*}$ $3$ $3\omega^{2}$ $3\omega$ 00 00 -$\text{-}1$$\text{-}\omega^{2}$$\text{-}\omega$$\text{-}1$$\text{-}\omega^{2}$$\text{-}\omega$$\text{-}1$$\text{-}\omega^{2}$$\text{-}\omega$
$\mathbf{3}^{(3)*}$ $3$ $3\omega^{2}$ $3\omega$ 00 00 $1$ $\omega^{2}$ $\omega$ $i$ $i\omega^{2}$ $i\omega$$\text{-}i$$\text{-}i\omega^{2}$$\text{-}i\omega$
$\mathbf{4}$ $4$ $4$ $4$ $1$$\text{-}2$ 00 00 00 00 00 00 00 00 00
$\mathbf{4}^{\prime}$ $4$ $4$ $4$$\text{-}2$ $1$ 00 00 00 00 00 00 00 00 00
III Basic Gates
In this work, we consider gate sets for both qubit and hybrid qubit-qutrit systems. Our qubit decompositions use the well-known fault tolerant Clifford + T gate set [104]. This choice is informed by the expectation that quantum simulations for lattice gauge theories will ultimately require fault tolerance to achieve quantum advantage [105, 106, 100, 101, 23]. Throughout, we adopt the notation $\oplus_{m}$ to mean addition mod $m$.
For conciseness, we use a larger than necessary gate set which we will later decomposes in terms of T-gate to obtain resource costs. The single qubit gates used are the Pauli rotations, $R_{\alpha}(\theta)=e^{i\theta\alpha/2}$, where $\alpha=X,Y,Z$. We also consider four entangling operations: SWAP, CNOT, and multicontrolled CnNOT, and the controlled SWAP (CSWAP). The two-qubit operations can be written as
$\text{SWAP}~{}|a\rangle|b\rangle=|b\rangle|a\rangle,$ (8)
$\text{CNOT}~{}|a\rangle|b\rangle=|a\rangle|b\oplus_{2}a\rangle.$ (9)
while the multicontrolled generalizations are
$\text{C}^{n}\text{NOT}~{}\Big{(}\prod_{n}|a_{n}\rangle\Big{)}|b\rangle=\Big{(}\prod_{n}|a_{n}\rangle\Big{)}|b\oplus_{2}\prod_{n}a_{n}\rangle,$ (10)
$\text{CSWAP}|a\rangle|b\rangle|c\rangle=|a\rangle|b(1\oplus_{2}a)\oplus_{2}ac\rangle|c(1\oplus_{2}a)\oplus_{2}ba\rangle.$ (11)
The hybrid encoding uses a more novel set of single, double, and triple qudit gates. The single-qudit two level rotations we consider are denoted by $R^{\alpha}_{b,c}(\theta)$, where $\alpha=\{X,~{}Y,~{}Z\}$ and indicate a Pauli-style rotation between levels $b$ and $c$. The subscripts will be omitted to indicate that the operation is performed on a qubit rather than a qutrit state. Additionally, we account for the primitive two-qudit gate,
$C^{a}_{b}X^{c}_{d,e}~{},$ (12)
which corresponds to the CNOT operation controlled on state $b$ of qubit $a$, and targets qubit $c$ with an $X$ operation between the levels $d$ and $e$. We also for conciseness consider the CSum gate:
$\textsc{CSum}^{a,b}|i\rangle_{a}|j\rangle_{b}=|i\rangle_{a}|i\oplus_{3}j\rangle_{b}.$ (13)
which is a controlled operation on qubit or qutrit $a$ and targets qutrit $b$. It can be verified that that the CSum (see e.g. Ref. [10, 107]) gate is related to the $C^{a}_{b}X^{c}_{d,e}$ gates by
$\textsc{CSum}^{a,b}$ $=\left(C^{a}_{1}X^{b}_{0,1}C^{a}_{1}X^{b}_{1,2}\right)\left(C^{a}_{2}X^{b}_{1,2}C^{a}_{2}X^{b}_{0,1}\right).$ (14)
Finally, we consider multi-controlled versions of both of these gates. The gate $C^{a}_{b}C^{c}_{d}X^{e}_{f,g}$ corresponds to multicontrolled generalization of Eq. (12). The second multiqudit gate is the CCSum which acts as follows
$\textsc{CCSum}^{a,b,c}|i\rangle_{a}|j\rangle_{b}|k\rangle_{c}=|i\rangle_{a}|j\rangle_{b}|k\oplus_{3}ij\rangle_{c}.$ (15)
A unique artifact of this choice of quantum gates is that one can decompose multi-controlled gates using the traditional Toffoli staircase decomposition [108, 104, 109].
IV Primitive Gates
We present the primitive gates for a pure gauge theory in the following subsections using the methods developed in previous papers on the binary tetrahedral, $\mathbb{BT}$, and binary octahedral, $\mathbb{BO}$ groups [97, 98, 100]. Using this formulation confers at least two benefits: first, it is possible to design algorithms in a theory- and hardware-agnostic way; second, the circuit optimization is split into smaller, more manageable pieces. This construction begins with defining for a finite group $G$ a $G$-register by identifying each group element with a computational basis state |$\ket{g}$. Then, Ref. [97] showed that Hamiltonian time evolution can be performed using a set of primitive gates. These primitive gates are: inversion $\mathfrak{U}_{-1}$, multiplication $\mathfrak{U}_{\times}$, trace $\mathfrak{U}_{\rm Tr}$, and Fourier transform $\mathfrak{U}_{F}$ [97].
The inversion gate, $\mathfrak{U}_{-1}$, is a single register gate that takes a group element to its inverse:
$\mathfrak{U}_{-1}\ket{g}=\ket{g^{-1}}\text{.}$ (16)
The group multiplication gate acts on two $G-$registers. It takes the target $G-$register and changes the state to the left-product with the control $G-$register:
[…]
Left multiplication is sufficient for a minimal set as right multiplication can be implemented using two applications of $\mathfrak{U}_{-1}$ and $\mathfrak{U}_{\times}$, albeit optimal algorithms may take advantage of an explicit construction [96].
The trace of products of group elements appears in lattice Hamiltonians. We can implement these terms by combining $\mathfrak{U}_{\times}$ with a single-register trace gate:
$\mathfrak{U}_{\operatorname{Tr}}(\theta)\left|g\right>=e^{i\theta\operatorname{Re}\operatorname{Tr}g}\left|g\right>.$ (18)
The final gate required is the group Fourier transform $\mathfrak{U}_{F}$. The Fourier transform of a finite $G$ is defined as
$\hat{f}(\rho)=\sqrt{\frac{d_{\rho}}{|G|}}\sum_{g\in G}f(g)\rho(g),$ (19)
where $|G|$ is the size of the group, $d_{\rho}$ is the dimensionality of the representation $\rho$, and $f$ is a function over $G$. The gate that performs this acts on a single $G$-register with some amplitudes $f(g)$ which rotate it into the Fourier basis:
$\mathfrak{U}_{F}\sum_{g\in G}f(g)\left|g\right>=\sum_{\rho\in\hat{G}}\hat{f}(\rho)_{ij}\left|\rho,i,j\right>.$ (20)
The second sum is taken over $\rho$, the irreducible representations of $G$; $\hat{f}$ denotes the Fourier transform of $f$. After performing the Fourier transform, the register is denoted as a $\hat{G}$-register to indicate the change of basis. A schematic example of this gate is show in Fig. 2.
While it is possible to pass the matrix constructed in Fig. 2 into a transpiler, more efficient methods for constructing these operators exist [110, 102, 111, 112, 113, 114]. While method varies in their actualization, the underlying spirit is the same as for the discrete quantum Fourier transformation. The principle method involves building the quantum Fourier transformation up through a series of subgroups. In [114], it was shown that instead of the exponential $\mathcal{O}(4^{n})$ scaling for traditional transpilation, the quantum Fourier transform scales like $\mathcal{O}(\text{polylog}(|2^{n}|))$, where polylog indicates a polynomial of logarithms.
Figure 2: Construction of $\mathfrak{U}_{F}$ from Eq. (19) using column vectors $\tilde{\rho}_{i,j}=\sqrt{d_{\rho}/|G|}\rho_{i,j}$ where $\rho_{i,j}=\rho_{i}(g_{j})$.
In the rest of this paper, we will construct each of these primitive gates, and evaluate the overall cost. For each gate, we will start with a pure qubit system. Then, we will consider a register with three qutrits and two qubits as suggested by the group presentation in Eq. 1.
IV.1 Inversion Gate
Figure 3: T-gate optimized version of $\mathfrak{U}_{-1}$ for $\Sigma(36\times 3)$. The letter indicates the generator and subscript indicates the qubit in the generator register. This implementation requires 119 T-gates and 4 ancilla.
Figure 4: $\mathfrak{U}_{-1}$, for $\Sigma(36\times 3)$ using qutrit-qubit encoding.
For the construction of $\mathfrak{U}_{-1}$, we first write the inverse of the group element $g$ as
$g^{-1}=\omega^{2p}V^{t+2t}V^{2s}E^{2r}C^{2q}=\omega^{p^{\prime}}C^{q^{\prime}}E^{r^{\prime}}V^{2s^{\prime}+t^{\prime}}.$ (21)
where the permutations rules are found to be
$p^{\prime}=$ $2p\oplus_{3}qr\oplus_{3}2qrt,$
$q^{\prime}=$ $2(q\oplus_{3}qs\oplus_{3}2qt\oplus_{3}rt\oplus_{3}2qst\oplus_{3}rst),$
$r^{\prime}=$ $2(r\oplus_{3}rs\oplus_{3}2qt\oplus_{3}2rt\oplus_{3}2qst\oplus_{3}2rst),$
$s^{\prime}=$ $s\oplus_{2}t,$
$t^{\prime}=$ $t.$ (22)
A detailed derivation of the permutation rules and the associated $\mathfrak{U}_{-1}$ is found in App. A along with two other forms of $\mathfrak{U}_{-1}$ which use fewer ancilla. The idealized qubit circuit is shown in Fig. 3 and requires 119 T-gates and 4 clean ancilla. The qubit-qutrit hybrid encoding $\mathfrak{U}_{-1}$ is found in Fig. 4.
IV.2 Multiplication Gate
The multiplication gate $\mathfrak{U}_{\times}$ takes two $G$-registers storing two group elements $g=\omega^{p}C^{q}E^{r}V^{2s+t}$ and $h=\omega^{p^{\prime}}C^{q^{\prime}}E^{r^{\prime}}V^{2s^{\prime}+t^{\prime}}$ and stores into the second register the group element $gh=\omega^{p^{\prime\prime}}C^{q^{\prime\prime}}E^{r^{\prime\prime}}V^{2s^{\prime\prime}+t^{\prime\prime}}$. Using the reordering relations of Eq. (7) one can derive that
$p^{\prime\prime}=$ $p\oplus_{3}p^{\prime}\oplus_{3}q^{\prime}r\oplus_{3}q^{\prime}rs\oplus_{3}2q^{\prime}rt\oplus_{3}2q^{\prime}r^{\prime}t$
$\oplus_{3}2rr^{\prime}t\oplus_{3}2q^{\prime}rst\oplus_{3}2rr^{\prime}st,$
$q^{\prime\prime}=$ $q+q^{\prime}+sq^{\prime}\oplus_{3}2tq^{\prime}\oplus_{3}2stq^{\prime}\oplus_{3}2tr^{\prime}\oplus_{3}2str^{\prime},$
$r^{\prime\prime}=$ $r\oplus_{3}tq^{\prime}\oplus_{3}stq^{\prime}\oplus_{3}r^{\prime}\oplus_{3}sr^{\prime}\oplus_{3}2tr^{\prime}\oplus_{3}2str^{\prime},$
$s^{\prime\prime}=$ $(s\oplus_{2}s^{\prime}\oplus_{2}tt^{\prime}),$
$t^{\prime\prime}=$ $(t\oplus_{2}t^{\prime}).$ (23)
These rules are rather clunky and in order to write a systematic multiplication gate we decompose $\mathfrak{U}_{\times}$ into the following product,
$\mathfrak{U}_{\times}=\mathfrak{U}_{\times,\omega}\mathfrak{U}_{\times,C}\mathfrak{U}_{\times,E}\mathfrak{U}_{\times,V^{2}}\mathfrak{U}_{\times,V},$ (24)
where $\mathfrak{U}_{\times,O}$ indicate multiplying the state of the $O$ generator register from the $g_{1}$ register onto the $g_{2}$ register. The breakdown of using this method and the product rules from Eq. (IV.2) yields the circuits composed in Fig. 5 and 6 for the two encodings.
Figure 5: Qubit implementation of $\mathfrak{U}_{\times}$ using the permutation gate $\chi$ and its inverse (both shaded orange) using 2 ancilla and has cost of 308 T-gates.
Figure 6: Qutrit+Qubit implementation of $\mathfrak{U}_{\times}$ using the permutation gate $\chi=X^{(0,1)}X^{(1,2)}$ and its inverse (both shaded orange).
IV.3 Trace Gate
There are two principle methods one could derive $\mathfrak{U}_{Tr}$. One method is to define a Hamiltonian of the form:
$\hat{H}_{Tr}=\sum_{g}Tr(g)|g\rangle\langle g|.$ (25)
Then, the trace operator can be written as $\mathfrak{U}_{Tr}\left(\theta\right)=\exp{\left(-\mathrm{i}\,\theta\,H_{{\rm Tr}}\right)}$. To obtain the matrix form of $\hat{H}_{Tr}$, one may fix a basis $\left\{|g_{1}\rangle,{...},|g_{|G|}\rangle\right\}$ where $|G|=108$ is the order of the group. In this basis, $\hat{H}_{Tr}$ is diagonal, and each diagonal entry is given by $H_{i,i}=Tr(g_{i})$.
To obtain a quantum circuit realizing $\mathfrak{U}_{{\rm Tr}}$, we use the tree-traversal algorithm developed in [115] [^3] The codes accompanying the cited publication can be found on this Github repository. which was shown to yield an exact circuit with asymptotically optimal CNOT gates count. The circuit obtained has 130 CNOT gates and 111 $R_{z}$ gates. Additional methods are also found in Ref. [116]. A second method for deriving this gate involves mapping group elements to their respective trace classes. $\Sigma(36\times 3)$ has 14 conjugacy classes that map to 10 different trace classes. If we only require the real part of the trace then this grouping reduces the 10 trace classes to 7 trace classes. The seven valid traces are ${\rm ReTr}(g)=\{3,-\frac{3}{2},0,\pm 1,\pm\frac{1}{2}\}$, which we can be labelled using three bits $\left(v_{0},v_{1},v_{2}\right)$ as shown in Table 3.
$\operatorname{Re}\operatorname{Tr}(g)$ $-3$ $-\frac{3}{2}$ $-1$ $-\frac{1}{2}$ 0 1 $\frac{1}{2}$
$v_{0}$ 0 1 0 1 0 0 1
$v_{1}$ 0 0 1 0 1 0 1
$v_{2}$ 0 0 1 1 0 1 1
Table 3: Map $(p,q,r,s,t)\mapsto(v_{0},v_{1},v_{2})$ via ${\rm{Re}Tr}(g)$.
This map can be represented as three boolean functions, one for each of the variables $v_{0}$, $v_{1}$ and $v_{2}$. For quantum computation, it is convenient to write boolean functions in the so-called exclusive-or sum of products (ESOP) form [117, 118]. Then, the function can be mapped to a quantum circuit in a straightforward manner since each term in the ESOPs corresponds to a Toffoli gate. For each function, we start with their minterm forms [119]. Then, we use the exorcism algorithm to find a simpler ESOP expression for each of the three functions. After factorizations, we show the final expressions in Eq. 26.
$v_{0}=$ $\bar{p}_{0}\left[1\oplus_{2}\bar{t}\left(q_{0}r_{1}\oplus_{2}q_{1}r_{0}\right)\oplus_{2}\bar{s}\left(\bar{q}_{0}\bar{r}_{0}\oplus_{2}\bar{q}_{1}\bar{r}_{1}\right)\right]$
$\oplus_{2}\bar{p}_{1}\left[t\oplus_{2}s\left(\bar{q}_{1}\bar{r}_{0}\oplus_{2}\bar{q}_{0}\bar{r}_{1}\right)\right]$
$\oplus_{2}\bar{p}_{1}\bar{t}(q_{0}\oplus_{2}\bar{q}_{1})(r_{0}\oplus_{2}\bar{r}_{1}),$
$v_{1}=$ $\bar{t}\oplus_{2}(q_{0}\oplus_{2}\bar{q}_{1})(r_{0}\oplus_{2}\bar{r}_{1})\bar{s}\bar{t},$
$v_{2}=$ $s\oplus\bar{s}t.$ (26)
$\mathfrak{U}_{Tr}$ can be decomposed as $\mathfrak{U}_{Tr}(\theta)=V\,\mathfrak{U}^{\prime}_{Tr}(\theta)V^{\dagger}$ where $V$ is a unitary operator realizing the map $\left(p,q,r,s,t\right)\mapsto\left(v_{0},v_{1},v_{2}\right)$. This yields $\mathfrak{U}^{\prime}_{Tr}(\theta)\equiv e^{\mathrm{i}\theta H^{\prime}}$, where
$H^{\prime}=$ $\frac{3}{16}III+\frac{3}{16}IIZ+\frac{3}{16}IZI+\frac{5}{16}IZZ$
$+\frac{1}{16}ZII+\frac{9}{16}ZIZ+\frac{9}{16}ZZI+\frac{15}{16}ZZZ.$ (27)
Figure 7 shows the quantum circuit of the operator $V$ realizing the map $\left(p,q,r,s,t\right)\mapsto\left(v_{0},v_{1},v_{2}\right)$. Finally, the circuit of $\mathfrak{U_{Tr}}^{\prime}(\theta)$ is shown in Figure 8.
Figure 7: Quantum Circuit of the map $\left(p,q,r,s,t\right)\mapsto\left(v_{0},v_{1},v_{2}\right)$ from the group to the seven real trace classes ${\rm ReTr(g)}=\{3,-1.5,0,\pm 1,\pm 0.5\}$. This requires 15 Toffoli gates and thus 105 T gates
Figure 8: Quantum circuit of the operator $\mathfrak{U}^{\prime}_{Tr}(\theta)$ where we have set $\phi\equiv-\frac{\theta}{8}$.
Figure 9: Quantum Circuit of the map $\left(p,q,r,s,t\right)\mapsto\left(v_{0},v_{1},v_{2}\right)$ from the group to the seven real trace classes ${\rm ReTr(g)}=\{3,-1.5,0,\pm 1,\pm 0.5\}$ for qutrit+qubit. In analogy to the closed and open circle notation for control qubits, the black and white qutrit controls represent controlled-on or controlled-on-others e.g. a white 1 is a control qutrit which applies a gate if in the state |$\ket{0}$ or |$\ket{2}$.
IV.4 Fourier Transform Gate
The standard $n$-qubit quantum Fourier transform (QFT) [120] corresponds to the quantum version of the fast Fourier transform of $\mathbb{Z}_{2^{n}}$. Quantum Fourier transforms, $\mathfrak{U}_{QFT}$, over some nonabelian groups are known [111, 112, 102, 114, 98]. However, for all the crystal-like subgroups of interest to high energy physics $\mathfrak{U}_{QFT}$ is currently unknown [121] and there is not a clear algorithmic way to construct $\mathfrak{U}_{QFT}$ in general. Therefore, we instead construct a suboptimal $\mathfrak{U}_{F}$ from Eq. (19) using the irreps of Sec. II. The structure of $\mathfrak{U}_{F}$ is ordered as follows. The columns index $|g\rangle$ from |$\ket{0}$ to |$\ket{256}$ according to Eq. (1). We then index the irreducible representation $\rho_{i}$ ordered sequentially from $i=1$ to $i=14$.
Since $\Sigma(36\times 3)$ has 108 elements, on a qubit device $\mathfrak{U}_{F}$ must be embedded into a larger $2^{8}\times 2^{8}=256\times 256$ matrix [^4] An explicit construction of the matrix $\mathfrak{U}_{F}$ is found in the supplementary information. While the matrix This matrix was then passed to the Qiskit v0.43.1 transpiler, and an optimized version of $\mathfrak{U}_{F}$ needed 30956 CNOTs, 2666 $R_{X}$, 32806 $R_{Y}$, and 55234 $R_{Z}$ gates; the Fourier gate is the most expensive qubit primitive. As will be discussed in Sec. V, $\mathfrak{U}_{F}$ dominates the total simulation costs and future work should be devoted to finding a $\Sigma(36\times 3)$ $\mathfrak{U}_{QFT}$
For the hybrid qubit-qutrit implementation, $\mathfrak{U}_{F}$ is of dimensions $108\times 108$. To obtain a quantum circuit, we built a qubit-qutrit compiler, see Ref. [122]. The outline of the compilation is discussed in Appendix B. The compiler uses the gates discussed in Sec. III, and Table 4 shows its resulting gates count.
Basic Gate $\mathcal{U}_{-1}$ $\mathcal{U}_{\times}$ $\mathcal{U}_{Tr}$ $\mathcal{U}_{F}$
$R_{\alpha}$ 0 0 7 92,568
$R^{\alpha}_{b,c}$ 3 0 0 35,310
CNOT 1 2 9 16,656
C2NOT 0 1 2 0
C3NOT 0 0 0 0
C4NOT 0 0 0 0
$C_{2}X_{b,c}$ 2 3 0 67,260
$C_{3}X_{b,c}$ 12 20 1 8,988
$C_{3}C_{2}X_{b,c}$ 8 12 13 0
$C_{3}C_{3}X_{b,c}$ 4 36 45 0
Table 4: Gate cost of primitive gates for $\Sigma(36\times 3)$ for a qutrit-qubit architecture. The costs for $\mathcal{U}_{F}$ were obtained with the hybrid compiler described in Appendix B. $C_{d}X_{b,c}$ refers to an $X$ rotation between states $b,c$ of a qutrit controlled by a qudit with $d$ levels
V Resource Costs
The relatively deep circuits presented above strongly suggest that simulating $\Sigma(36\times 3)$ will require error correction. The preclusion of universal transversal sets of gates stated in the Eastin-Knill theorem [123] requires compromises be made. In most error correcting codes, the Clifford gates are designed to be transversal [104, 124, 125, 126, 127]. This leaves the nontransversal T gate as the dominant cost of fault-tolerant algorithms [104, 128]. Beyond these standard codes, novel universal sets exist with transversal $\mathbb{BT},\mathbb{BO},\mathbb{BI}$ and He(3) [^5] The Heisenberg group of dimension 3, which is a noncrystal-like subgroup of SU(3) gates [129, 130, 131, 132, 133] which warrant exploration for use in lattice gauge theory.
Table 5: Number of physical T gates and clean ancilla required to implement logical gates for (top) basic gates taken from [104] (bottom) primitive gates for $\Sigma(36\times 3)$.
Gate T gates Clean ancilla
C2NOT 7 0
C3NOT 21 1
C4NOT 35 2
CSWAP 7 0
$R_{Z}$ 1.15 $\log_{2}(1/\epsilon)$ 0
$\mathcal{U}_{-1}$ 119 4
$\mathcal{U}_{\times}$ 308 2
$\mathcal{U}_{Tr}$ $210+8.05\log_{2}(1/\epsilon)$ 7
$\mathcal{U}_{F}$ $185897.5\log_{2}(1/\epsilon)$ 0
For this work we will consider the following decompositions of gates into T gates for our resource estimates. First, while the CNOT is transversal, the Toffoli gate decomposes into six CNOTs and seven T gates [104]. With this, one can construct any CnNOT gates using $12\lceil\log_{2}{n}\rceil-$ 1 Toffoli gates and $n-2$ dirty ancilla qubits [^6] A clean ancilla is in state |$\ket{0}$. Dirty ancillae have unknown states. which can be reused later [108, 104, 109]. For the $R_{Z}$ gates, we use the repeat-until-success method of [134] which finds these gates can be approximated to precision $\epsilon$ with on average $1.15\log_{2}(1/\epsilon))$ T gates (and at worst $-9+4\log_{2}(1/\epsilon)$ [135]). For $R_{Y}$ and $R_{X}$, one can construct them with at most 3 $R_{Z}$. Putting all together, we can construct gate estimates for $\Sigma(36\times 3)$ (See Tab. 5).
Table 6: Number of primitive gates per link per $\delta t$ neglecting boundary effects as a function of $d$ for $H_{KS}$ and $H_{I}$.
Gate $\mathfrak{U}_{F}$ $\mathfrak{U}_{\rm Tr}$ $\mathfrak{U}_{-1}$ $\mathfrak{U}_{\times}$
$e^{-i\delta H_{KS}}$ 2 $\frac{1}{2}(d-1)$ $3(d-1)$ $6(d-1)$
$e^{-i\delta H_{I}}$ 4 $\frac{3}{2}(d-1)$ $2+11(d-1)$ $4+26(d-1)$
While the results in Tab. 5 are nearly optimal for the $\mathfrak{U}_{\operatorname{Tr}}$, $\mathfrak{U}_{\times}$, and $\mathfrak{U}_{-1}$, the result for $\mathfrak{U}_{F}$ is not. In Ref. [110] the authors show explicit demonstrations of an efficient decomposition of the nonabelian quantum Fourier transform $\mathfrak{U}_{QFT}$ using the methods of [102] for certain $SU(2)$ and $SU(3)$ subgroups. Since it is expected that the gate cost for Fourier transforms should scale as a polynomial of logarithms of the group size [114], one can perform a fit from the results in Ref. [110] to obtain an order of magnitude estimate for $\mathfrak{U}_{QFT}$ of $147+75\rm{log}_{2}(1/\epsilon)$ – a factor of $\sim 2000$ smaller than our $\mathfrak{U}_{F}$.
Clearly, the cost of simulating $\Sigma(36\times 3)$ depends on $\epsilon$. To optimize the cost, the synthesis error from finite $\epsilon$ should be balanced with other sources of error in the quantum simulation like Trotter error, discretization error, and finite volume error. These other sources of error are highly problem-dependent, but here we will follow prior works [136, 42, 137] and take a fiducial $\epsilon=10^{-8}$.
Primitive gate costs for implementing $H_{KS}$ [138] and $H_{I}$ [96], per link per Trotter step $\delta t$ are shown in Tab. 6. Using this result, we can determine the total T gate count $N^{H}_{T}=C^{H}_{T}\times dL^{d}N_{t}$ for a $d$ spatial lattice simulated for a time $t=N_{t}\delta t$. We find that for $H_{KS}$
$C^{KS}_{T}=2310(d-1)+(371791+4.025d)\log_{2}\frac{1}{\epsilon}.$ (28)
With this, the total synthesis error $\epsilon_{T}$ can be estimated as the sum of $\epsilon$ from each $R_{Z}$. In the case of $H_{KS}$ this is
$\epsilon_{T}=\frac{1}{2}(646593+7d)dL^{d}N_{t}\times\epsilon.$ (29)
If one looks to reduce lattice spacing errors for a fixed number of qubits, one can use $H_{I}$ which would require
$C^{I}_{T}=9632d-8192+(743578+12.075d)\log_{2}\frac{1}{\epsilon},$ (30)
where the total synthesis error is
$\epsilon_{T}=\frac{1}{2}(1293179+21d)dL^{d}N_{t}\times\epsilon.$ (31)
Following [139, 136, 100], we will make resource estimates based on our primitive gates for the calculation of the shear viscosity $\eta$ on a $L^{3}=10^{3}$ lattice evolved for $N_{t}=50$, and total synthesis error of $\epsilon_{T}=10^{-8}$. Considering only the time evolution and neglecting state preparation (which can be substantial [140, 141, 142, 143, 144, 145, 146, 147, 75, 148, 149, 150, 151, 150, 152, 153, 154]), Kan and Nam estimated $6.5\times 10^{48}$ T gates would be required for an pure-gauge $SU(3)$ simulation of $H_{KS}$. This estimate used a truncated electric-field digitization and considerable fixed-point arithmetic – greatly inflating the T gate cost. Here, using $\Sigma(36\times 3)$ to approximate $SU(3)$ requires $7.0\times 10^{12}$ T gates for $H_{I}$ and $3.5\times 10^{12}$ T gates for $H_{KS}$. The T gate density is roughly 1 per $\Sigma(36\times 3)-$register per clock cycle. Thus $\Sigma(36\times 3)$ reduces the gate costs of [136] by $10^{36}$. Similar to the previous results for discrete groups of SU(2), $\mathfrak{U}_{F}$ dominates the simulations – being over 99% of the computation regardless of Hamiltonian. However [110] shows that the Fourier transformation for $\mathbb{BT}$ and $\mathbb{BO}$ can be brought down. Using the estimate for $\Sigma(36\times 3)$, the Fourier gate contribution is reduced to only 51% of the simulation with a reduced total T gate count of $5.7\times 10^{9}$ for $H_{I}$ with $L=10$
VI Outlook
This article provided a construction of primitive gates necessary to simulate a pure SU(3) gauge theory via a discrete subgroup $\Sigma(36\times 3)$. In addition, we have also estimated the T-gate cost incurred to compute the shear viscosity using the $\Sigma(36\times 3)$ group. Notably, we found that our construction improves the T-gate cost upon that of Ref. [136] by 36 orders of magnitude. This cost reduction comes at the expense of model accuracy.
For both qubit and hybrid qubit-qutrit implementations, $\mathfrak{U}_{F}$ dominates the cost suggesting that further reductions by identifying a $\mathfrak{U}_{QFT}$ for $\Sigma(36\times 3)$. In fact, as demonstrated in Ref. [110], the cost of a $\mathfrak{U}_{QFT}$ versus $\mathfrak{U}_{F}$ can be as large as a factor of $\sim 2000$.
In addition, the much-improved overall cost due to the use of the $\Sigma(36\times 3)$ group supports the need to also study other discrete subgroups of SU(2) and SU(3). To this end, recent studies (e.g. Ref [100, 101]) have already constructed primitive gates for some SU(2) discrete subgroups, the binary tetrahedral and binary octahedral. It remains to develop such gates for other subgroups, for example the larger subgroups of SU(3) such as $\Sigma(72\times 3)$, $\Sigma(216\times 3)$ and $\Sigma(360\times 3)$ as well as the $\mathbb{BI}$ group. The larger groups will reduce discretization errors but at the cost of a longer circuit depth.
Finally, beyond pure gauge, approximating QCD requires incorporating fermion fields [155, 79, 156]. Many methods exist to incorporate staggered and Wilson fermions. It is worth comparing the resource costs for explicit spacetime simulations using staggered versus Wilson fermions not only in terms of T-gates but also spacetime costs using methods such as [157].
Acknowledgements.
The authors would like to thank Sohaib Alam and Stuart Hadfield for helpful feedback. This material is based on work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. EG was supported by the NASA Academic Mission Services, Contract No. NNA16BD14C and NASA-DOE interagency agreement SAA2-403602. YJ is grateful for the support of Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant SFB TR 110/2. EM is supported by the U.S. Department of Energy, Office of Nuclear Physics under Award Numbers DE-SC021143 and DE- FG02-95ER40907. SZ is supported by the National Science Foundation CAREER award (grant CCF-1845125). Part of this research was performed while SZ was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). Fermilab is operated by Fermi Research Alliance, LLC under contract number DE-AC02-07CH11359 with the United States Department of Energy.
Appendix A Derivation of Inversion Gates
The inversion rules from Eq. (IV.1) are written in a three level notation. If one wants to simulate systems using qubits one needs to map these to a qudit basis rules to qubit basis rules. We first begin with the $p^{\prime}$ rule
$p^{\prime}=$ $2p\oplus_{3}qr(1-t)$
$=$ $2(p_{0}\oplus_{3}2p_{1})\oplus_{3}(q_{0}\oplus_{3}2q_{1})(r_{0}\oplus_{3}2r_{1})(1-t)$
$=$ $2(p_{0}\oplus_{3}2p_{1})\oplus_{3}(1-t)\big{(}q_{0}r_{0}\oplus_{3}2q_{1}r_{0}$
$\oplus_{3}2r_{1}q_{0}\oplus_{3}4q_{1}r_{1}\big{)}.$
In order to turn this trinary arithmetic into binary arithmetic we need the following transformation axiom:
$p^{\prime}=$ $p\oplus_{3}1,$
$p_{0}^{\prime}=$ $p_{0}\oplus_{2}p_{1}\oplus_{2}1,$
$p_{1}^{\prime}=$ $p_{0},$ (32)
and
$p^{\prime}=$ $p\oplus_{3}2,$
$p_{0}^{\prime}=$ $p_{1},$
$p_{1}^{\prime}=$ $p_{1}\oplus_{2}p_{0}\oplus_{2}1.$ (33)
Using this set of transformation rules we find
$p_{0}^{\prime}=$ $p_{1}\oplus_{2}(1\oplus_{2}t)\big{(}(1\oplus_{2}p_{0})(q_{0}r_{0}\oplus_{2}q_{1}r_{1})$
$\oplus_{2}p_{0}(q_{1}r_{0}\oplus_{2}r_{1}q_{0})\big{)},$
$p_{1}^{\prime}=$ $p_{0}\oplus_{2}(1\oplus_{2}t)\big{(}(1\oplus_{2}p_{1})(q_{1}r_{0}\oplus_{2}q_{0}r_{1})$
$\oplus_{2}p_{1}(q_{0}r_{0}\oplus_{2}q_{1}r_{1})\big{)},$
$q_{0}^{\prime}=$ $q_{1}(1\oplus_{2}s)(1\oplus_{2}t)\oplus_{2}q_{0}s(1\oplus_{2}t)$
$\oplus_{2}r_{1}(1\oplus_{2}s)t\oplus_{2}r_{0}st,$
$q_{1}^{\prime}=$ $q_{0}(1\oplus_{2}s)(1\oplus_{2}t)\oplus_{2}q_{1}s(1\oplus_{2}t)$
$\oplus_{2}r_{0}(1\oplus_{2}s)t\oplus r_{1}st,$
$r_{0}^{\prime}=$ $r_{1}(1\oplus_{2}s)(1\oplus_{2}t)\oplus_{2}r_{0}s(1\oplus_{2}t)$
$\oplus q_{0}t(1\oplus_{2}s)\oplus q_{1}ts,$
$r_{1}^{\prime}=$ $r_{0}(1\oplus_{2}s)(1\oplus_{2}t)\oplus_{2}r_{1}s(1\oplus_{2}t)$
$\oplus_{2}q_{1}t(1\oplus_{2}s)\oplus_{2}q_{0}ts.$ (34)
Naively translating these rules as written yields the circuit provided in Fig. 10. However the resource cost of 420 T-gates can be optimized significantly. By clever use of ancillae one could reduce the T-gate costs down to 203 T-gates using the circuit provided in Fig. 11.
Figure 10: $\mathfrak{U}_{-1}$ for $\Sigma(108)$ which requires 420-T gates.
Instead of writing a circuit for the circuit for the whole inversion ruleset of Eq. (IV.1), one instead could use commutation rules to reduce the T-gate costs even further. This construction allows the inversion operation to be decomposed into a product of smaller operations:
$\mathfrak{U}_{-1}=\mathfrak{U}_{-1}^{C}\mathfrak{U}_{-1}^{E}\mathfrak{U}_{-1}^{V^{2}}\mathfrak{U}_{-1}^{V}\mathfrak{U}_{-1}^{l}.$ (35)
$\mathfrak{U}_{-1}^{l}$ takes each local generator to its inverse:
$t\rightarrow t$
$s\rightarrow s\oplus_{2}t$
$r\rightarrow 2r$
$q\rightarrow 2q$
$p\rightarrow 2p.$
The operation $\mathfrak{U}_{-1}^{V}$ involves propagating through the operator $V^{t}$ until it is the right most element. This yields the transformations:
$s\rightarrow s$
$r\rightarrow r(1-t)\oplus_{3}qt$
$q\rightarrow 2rt\oplus_{3}q(t-1)$
$p\rightarrow p\oplus_{3}rq(1-t).$
The generators $C$ and $E$ are normal ordered at this point. The operation $\mathfrak{U}_{-1}^{V^{2}}$ has the following transformation rule
$r\rightarrow 2rs\oplus_{3}r(1-s)$
$q\rightarrow 2qs\oplus_{3}q(1-s)$
$p\rightarrow p.$
At this point the transformation rules for $C,~{}E,$ and $\omega$ are trivial. After all these suboperations are constructed we end up with the inversion operation from the main text provided in Fig. 3 and Fig. 4.
Figure 11: A more $T$-gate optimized version of $\mathfrak{U}_{-1}$ for $\Sigma(108)$ which requires requires 203-T gates and 2 ancilla.
Appendix B Qubit-Qutrit Compiler
This section details the compilation method, and the full codes can be found in Ref. [122]. The overarching approach of this compiler is to generalize the qubit Quantum Shannon Decomposition (QSD) [158, 159] to apply to a register with qubits and qutrits. We will consider a unitary operator $U$ acting on $n_{1}$ qubits and $n_{2}$ qutrits; that is $U$ is of dimension $N\times N$ where $N=2^{n_{1}}\times 3^{n_{2}}$.
First, let’s organize the qudits as $q_{1},q_{2},{...},q_{n_{1}+n_{2}}$. We will say that the left-most qudit is the top qudit. The compiler iteratively performs the qubit QSD if the top qudit is a qubit, and otherwise performs our realization of the qutrit QSD. The process eventually terminates when we reach the bottom qudit, in which case, we use either a single-qubit gate decomposition or a single-qutrit gate decomposition depending on whether the bottom qudit is qubit of qutrit. For the single qubit gate, we use the Euler angle parametrization $ZYZ$ and for the single qutrit gate, we use the decomposition given in Ref. [107].
It is convenient to start with the qubit QSD case. In this case, a Cosine-Sine decomposition (CSD) (see e.g. Refs. [160, 161, 162]) is first performed, resulting in
$U=(V_{1}\oplus V_{2})\begin{pmatrix}C&-S\\ S&C\\ \end{pmatrix}(W_{1}\oplus W_{2}),$ (36)
where $V_{1,2}$, $W_{1,2}$ are unitaries with dimension $N/2$. $C$ and $S$ are diagonal matrices e.g $C=\rm{diag}\left(\cos\theta_{1},{...},\cos\theta_{N/2}\right)$ and similarly for $S$.
Following [158], the next step is to decompose the two block-diagonal unitary matrices:
$\left(V_{1}\oplus V_{2}\right)$ $=(\mathbb{1}\otimes M)(D\oplus D^{\dagger})(\mathbb{1}\otimes N).$ (37)
where $M$ and $N$ are unitaries acting only on $n_{1}-1$ qubits and $n_{2}$ qutrits, and $D$ is a diagonal unitary of dimension $N/2$. Thus, the compilation problem is reduced to decomposing the unitaries $D\oplus D^{\dagger}$ and the $CS$ into simple gates. The $D\oplus D^{\dagger}$ can be implemented as a uniformly controlled rotations on the top qudit (see e.g. Ref. [158]). The $CS$ matrix on the other hand, is related to $D\oplus D^{\dagger}$ by the rotation gate $R_{x}(\pi/2)$ on the top qubit. This concludes the case the top qudit is a qubit.
In the case that the top qudit is qutrit, we need to find a qutrit realization of the procedure above. The starting point is to perform two CSD as in e.g. Ref. [163]. This decomposition reads as
$U=\left(V_{1}\oplus V_{2}\oplus V_{3}\right)$ $\left(\mathbb{1}\oplus D\oplus D^{\dagger}\right)\left(W_{1}\oplus W_{2}\oplus W_{3}\right)\begin{pmatrix}C&&-S\\ &\mathbb{1}&\\ S&&C\end{pmatrix}\left(V^{\prime}_{1}\oplus V^{\prime}_{2}\oplus V^{\prime}_{3}\right)\left(\mathbb{1}\oplus D^{\prime}\oplus D^{\prime\dagger}\right)\left(W^{\prime}_{1}\oplus W^{\prime}_{2}\oplus W^{\prime}_{3}\right),$ (38)
where each block is a unitary of dimension $N/3$. The blocks $C$, $S$ $D$ and $D^{\prime}$ are defined analogously to the qubit case.
The rest is to decompose the remaining block diagonal unitaries. By performing the decomposition in Ref. [158] twice, we obtain the relation
$V_{1}\oplus V_{2}\oplus V_{3}=\left(\mathbb{1}_{3\times 3}\otimes M\right)\left(D\oplus D\oplus D^{\dagger}\right)\left(\mathbb{1}_{3\times 3}\otimes N\right)$ $\left(D^{\prime}\oplus D^{\prime\dagger}\oplus\mathbb{1}\right)\left(\mathbb{1}_{3\times 3}\otimes M^{\prime}\right)\left(D^{\prime\prime}\oplus D^{\prime\prime}\oplus D^{\prime\prime\dagger}\right)\left(\mathbb{1}_{3\times 3}\otimes N^{\prime}\right).$ (39)
The unitaries $\mathbb{1}\oplus D\oplus D^{\dagger}$, $D\oplus D^{\dagger}\oplus\mathbb{1}$, $D\oplus D\oplus D^{\dagger}$ and $CS$ are uniformly controlled rotations. We can focus on the diagonal blocks because as in the qubit case, the $CS$ matrix can be diagonalized with an appropriate $R^{i,j}_{x}(\pi/2)$ on the top qutrit. Ref. [164] outlines the decomposition of qutrits uniformly controlled rotations in terms of single- and two-qutrit gates. A generalization can be obtained simply by limiting a $C^{a}_{b}X^{c}_{i,j}$ gate to $C^{a}_{1}X^{c}_{i,j}$ when the control qudit is a qubit.
For $n=2$ qubits, there exists optimal compilation algorithms (see e.g. Ref. [165]). Therefore, when the bottom two qudits are both qubits, we stop the decomposition and use Qiskit transpiler to obtain a quantum circuit.